# YOLO-NAS-Bench: A Surrogate Benchmark with Self-Evolving Predictors for YOLO Architecture Search

### 10 Mar 2026 v1

@@ Abstract
Neural Architecture Search (NAS) for object detection is severely bottlenecked by high evaluation cost, as fully training each candidate YOLO architecture on COCO demands days of GPU time. Meanwhile, existing NAS benchmarks largely target image classification, leaving the detection community without a comparable benchmark for NAS evaluation. To address this gap, we introduce YOLO-NAS-Bench, the first surrogate benchmark tailored to YOLO-style detectors. YOLO-NAS-Bench defines a search space spanning channel width, block depth, and operator type across both backbone and neck, covering the core modules of YOLOv8 through YOLO12. We sample 1,000 architectures via random, stratified, and Latin Hypercube strategies, train them on COCO-mini, and build a LightGBM surrogate predictor. To sharpen the predictor in the high-performance regime most relevant to NAS, we propose a Self-Evolving Mechanism that progressively aligns the predictor’s training distribution with the high-performance frontier, by using the predictor itself to discover and evaluate informative architectures in each iteration. This method grows the pool to 1,500 architectures and raises the ensemble predictor’s $R^{2}$ from 0.770 to 0.815 and Sparse Kendall Tau from 0.694 to 0.752, demonstrating strong predictive accuracy and ranking consistency. Using the final predictor as the fitness function for evolutionary search, we discover architectures that surpass all official YOLOv8–YOLO12 baselines at comparable latency on COCO-mini, confirming the predictor’s discriminative power for top-performing detection architectures.
@@ 1 Introduction
  • We design a YOLO-oriented search space covering backbone and neck with channel, depth, and operator dimensions that span the key modules of YOLOv8–YOLO12. Through random, stratified, and Latin Hypercube sampling, we build a benchmark database of 1,000 architectures fully trained on COCO-mini.
@@ 3 Method
Figure 2: Overview of the YOLO-NAS-Bench pipeline. (1) A YOLO-style search space spanning channel, depth, and operator dimensions across both backbone and neck is defined. (2) 1,000 architectures are sampled via three complementary strategies and trained on COCO-mini. (3) A LightGBM predictor is trained on the resulting {architecture, mAP} pairs. (4) The Self-Evolving Predictor iteratively expands the pool with high-value architectures discovered by evolutionary search, and retrains the predictor over 10 rounds, yielding an ensemble of 10 LightGBM models over 1,500 architectures.
An overview of the YOLO-NAS-Bench is illustrated in Fig. 2. Starting from a carefully designed search space over YOLO-style architectures, we sample and fully train 1,000 architectures on COCO-mini to build a ground-truth performance database. Each configuration is converted into a compact feature vector, upon which a LightGBM surrogate predictor is trained. To further strengthen the predictor in the high-performance regime that matters most for NAS, we propose a Self-Evolving Predictor loop that iteratively discovers, trains, and assimilates promising architectures.
@@ 3.1 Benchmark Construction
  • Channel width: Each of the four backbone stages (P2–P5) has an independently selectable channel count. The candidate sets grow with the stage depth to reflect the natural widening of feature hierarchies in modern detectors.
@@
  • Operator type: This dimension subsumes both the feature-extraction module in each stage and the downsampling operator between stages. Feature-extraction candidates include C2f [11], C3k2 and C2PSA [12], and C2fCIB [22]—the core building blocks employed across YOLOv8 through YOLO12—spanning a spectrum from lightweight convolutions (C2f) to efficient re-parameterized blocks (C3k2, C2fCIB) and attention-augmented modules (C2PSA). Downsampling candidates include standard Conv and the stride-channel-decoupled SCDown [22].
    In the backbone, the operator palette widens progressively: P2/P3 stages choose between {C2f, C3k2}, P4 adds C2fCIB, and P5 further includes C2PSA, reflecting the increasing complexity budget at lower resolutions. In the neck, channel width and depth are fixed to avoid excessive growth of the search space; only the operator type is searchable. Feature blocks N1–N4 share a single global choice from {C2f, C3k2, C2fCIB}, and the two downsample blocks D1–D2 share a single choice from {Conv, SCDown}. Upsample layers use fixed nn.Upsample throughout the YOLO family and are not searchable.
@@
Channel P2 128, 192, 256 3
P3 256, 320, 384, 448, 512 5
P4111P4 channels must be at least P3 to avoid degenerate configurations. 384, 512, 640, 768 4
P5 768, 1024, 1280 3
Depth P2 1, 2 2
P3 2, 3, 4 3
P4 2, 3, 4 3
P5 2, 3 2
Operator (backbone) P2/P3 C2f, C3k2 2
P4 C2f, C3k2, C2fCIB 3
P5 C2f, C3k2, C2fCIB, C2PSA 4
Operator (neck) N1–N4 C2f, C3k2, C2fCIB 3
@@
Each architecture configuration is converted into a 24-dimensional feature vector. Channel widths and block depths are represented as scalar values, while operator choices are one-hot encoded, allowing the tree-based predictor to split on each operator independently without imposing artificial ordering.
@@ 3.2 Self-Evolving Predictor
Evolutionary expansion (single round). Within each latency bucket, a target latency is sampled uniformly. An evolutionary algorithm (EA) [19] is then run with predicted mAP as the fitness function and real measured latency as the constraint. The EA uses a population of 50 and runs for 100 generations, adopts top 25% as parents, with offspring generated by uniform crossover (50%) and mutation (50%, mutation probability 0.2), and preserves the top 10% as elite. The final top 5 architectures from each bucket are retained on COCO-mini, yielding 50 new candidate architectures per round.
Iterative refinement. Each batch of 50 new architectures is fully trained on COCO-mini under the same protocol, their configuration and ground-truth mAP are recorded, and the pairs are merged into the architecture pool. The predictor is then retrained on the expanded pool. This loop repeats for 10 rounds, growing the pool from 1,000 to 1,500 architectures. Crucially, each round’s newly added architectures are biased toward the high-performance regions identified by the current predictor, progressively reducing the gap between the predictor’s training distribution and the distribution of architectures encountered during actual NAS search.
@@ 4.2 Main Results
Tab. 3 summarizes the surrogate predictor’s accuracy before and after Self-Evolving. The 10-model LightGBM ensemble trained on the initial 1,000 architectures achieves a validation sKT of 0.694. After 10 rounds of Self-Evolving, both $R^{2}$ and sKT get improved: $R^{2}$ rises from 0.770 to 0.815 (+4.5%), and sKT from 0.694 to 0.752 (+5.8%). An sKT of 0.752 indicates strong ranking consistency, while the $R^{2}$ of 0.815 suggests that the predictor captures the majority of variance in architecture performance. These results show that our surrogate predictor is a high-fidelity proxy for the true performance landscape, and confirm that the Self-Evolving mechanism effectively sharpens the predictor’s ranking ability. Fig. 4 further visualizes the close agreement between predicted and actual mAP across all 1,500 architectures.
@@
Large Medium-Large Medium Small
@@
YOLO12 32.9 54.13 30.8 42.51 30.4 22.17 27.2 20.40
YOLO11 32.8 37.73 32.1 29.84 30.8 20.03 27.7 15.30
YOLOv10 30.5 36.84 30.1 25.83 29.7 20.01 25.8 15.89
YOLOv9 32.3 53.07 31.6 26.85 31.6 25.59 27.1 19.25
YOLOv8 32.9 33.71 30.8 23.30 29.4 17.21 26.2 10.95
@@
As shown in Tab. 4, the four predictor-discovered architectures Pareto-dominate all official YOLO baselines (v8–v12) across the full latency spectrum, achieving consistently higher mAP at equal or lower latency. The advantage is most pronounced in the small-model regime, where Arch-D surpasses YOLO11s by +4.2% mAP at comparable latency, and remains substantial at the large end, where Arch-A exceeds YOLO12x in mAP while being $1.5{\times}$ faster. These results confirm that the predictor not only achieves high ranking correlation (sKT) on held-out data, but also exhibits strong discriminative power in the top-performance regime—architectures it deems promising indeed yield superior performance after full retraining.
@@ 4.3 Ablation Studies
To disentangle whether the predictor improvement stems from the Self-Evolving mechanism or merely from increasing the training pool size, we conduct an ablation: we train 200 randomly sampled architectures on COCO-mini, add them to the initial 1,000 architectures, and retrain the ensemble predictor. For fair comparison, we also truncate Self-Evolving when the pool reaches 1,200 architectures. Tab. 6 reports the results. Random expansion yields $R^{2}=0.776$ and sKT=0.701=0.701, while Self-Evolving achieves $R^{2}=0.798$ and sKT=0.738=0.738. The clear gap (+0.022 $R^{2}$, +0.037 sKT) confirms that the gain is attributable to the targeted enrichment of high-performance architectures by the Self-Evolving loop, rather than to the increase in data size alone.
@@
4.4 Limitations and future work.
The current benchmark is built on COCO-mini (10% of COCO) and measures latency on a single GPU (NVIDIA P40). To ensure the authenticity of latency measurements across different platforms, we do not model a latency predictor in the same way as mAP; thus, when using this benchmark for NAS, one still needs to measure the latency of each architecture empirically. But this cost is acceptable compared to the error that a latency predictor would introduce on unseen hardware. Extending YOLO-NAS-Bench to the full COCO dataset, diverse hardware platforms (edge GPUs, mobile NPUs), and additional tasks such as instance segmentation and pose estimation are natural next steps.
@@ 5 Conclusion
We present YOLO-NAS-Bench, the first surrogate benchmark tailored to YOLO-style detectors. By designing a search space that spans channel width, block depth, and operator type across backbone and neck, and by training 1,000 architectures sampled by different strategies on COCO-mini, we establish an initial performance database. Based on this, our Self-Evolving Predictor iteratively enriches the database with high-value architectures, raising the ensemble LightGBM predictor’s sKT from 0.694 to 0.752. Architectures discovered by predictor-guided evolutionary search surpass all official YOLOv8–YOLO12 baselines at comparable latency on COCO-mini, confirming the predictor’s practical utility.

### 18 May 2026 v2

@@ Abstract
Neural Architecture Search (NAS) for object detection is severely bottlenecked by high evaluation cost, as fully training each candidate YOLO architecture on COCO demands days of GPU time. Meanwhile, existing NAS benchmarks largely target image classification, leaving the detection community without a comparable benchmark for NAS evaluation. To address this gap, we introduce YOLO-NAS-Bench, the first surrogate benchmark tailored to YOLO-style detectors. YOLO-NAS-Bench defines a search space spanning channel width, block depth, and operator type across both backbone and neck, covering the core modules of YOLOv8 through YOLO12. We sample 1,000 architectures via random, stratified, and Latin Hypercube strategies, train them on COCO-mini, and build a LightGBM surrogate predictor. To sharpen the predictor in the high-performance regime most relevant to NAS, we propose a Self-Evolving Mechanism that progressively aligns the predictor’s training distribution with the high-performance frontier, by using the predictor itself to discover and evaluate informative architectures in each iteration. This method grows the pool to 1,500 architectures and raises the ensemble predictor’s $R^{2}$ from 0.770 to 0.815 and Sparse Kendall Tau from 0.694 to 0.752, demonstrating strong predictive accuracy and ranking consistency. Using the final predictor as the fitness function for evolutionary search, we discover architectures that surpass all official YOLOv8–YOLO12 baselines at comparable latency on COCO-mini, confirming the predictor’s discriminative power for top-performing detection architectures. The code is available at https://github.com/VDIGPKU/YOLO-NAS-Bench.
@@ 1 Introduction
  • We design a YOLO-oriented search space covering backbone and neck with channel, depth, and operator dimensions that span the key modules of YOLOv8–YOLO12. Through random, stratified, and Latin Hypercube sampling, we build an initial database of 1,000 architectures fully trained on COCO-mini.
@@ 3 Method
Figure 2: Overview of YOLO-NAS-Bench. (1) A YOLO-style search space spanning channel, depth, and operator dimensions across both backbone and neck is defined. (2) 1,000 architectures are sampled via three complementary strategies and trained on COCO-mini. (3) A LightGBM predictor is trained on the resulting {architecture, mAP} pairs. (4) The Self-Evolving Predictor iteratively expands the pool with high-value architectures discovered by evolutionary search, and retrains the predictor over 10 rounds, yielding an ensemble of 10 LightGBM models over 1,500 architectures.
An overview of YOLO-NAS-Bench is illustrated in Fig. 2. Starting from a carefully designed search space over YOLO-style architectures, we sample and fully train 1,000 architectures on COCO-mini to build a ground-truth performance database. Each configuration is converted into a compact feature vector, upon which a LightGBM surrogate predictor is trained. To further strengthen the predictor in the high-performance regime that matters most for NAS, we propose a Self-Evolving Predictor loop that iteratively discovers, trains, and assimilates promising architectures.
@@ 3.1 Benchmark Construction
  • Channel width: Each of the four backbone stages (S2–S5) has an independently selectable channel count. The candidate sets grow with the stage depth to reflect the natural widening of feature hierarchies in modern detectors.
@@
  • Operator type: This dimension subsumes both the feature-extraction module in each stage and the downsampling operator between stages. Feature-extraction candidates include C2f [11], C3k2 [12], C2fCIB [22], and A2C2f [20]—the core building blocks employed across YOLOv8 through YOLO12—spanning a spectrum from lightweight convolutions (C2f) to efficient re-parameterized blocks (C3k2, C2fCIB) and attention-augmented modules (A2C2f). Downsampling candidates include standard Conv and the stride-channel-decoupled SCDown [22].
In the backbone, the operator palette widens progressively: S2/S3 stages choose between {C2f, C3k2}, S4 adds C2fCIB, and S5 further includes A2C2f, reflecting the increasing complexity budget at lower resolutions. In the neck, channel width and depth are fixed to avoid excessive growth of the search space; only the operator type is searchable. Feature blocks N1–N4 share a single global choice from {C2f, C3k2, C2fCIB, A2C2f}, and the two downsample blocks D1–D2 share a single choice from {Conv, SCDown}. Upsample layers use fixed nn.Upsample throughout the YOLO family and are not searchable.
@@
Channel S2 128, 192, 256 3
S3 256, 320, 384, 448, 512 5
S4111S4 channels must be at least S3 to avoid degenerate configurations. 384, 512, 640, 768 4
S5 768, 1024, 1280 3
Depth S2 1, 2 2
S3 2, 3, 4 3
S4 2, 3, 4 3
S5 2, 3 2
Operator (backbone) S2/S3 C2f, C3k2 2
S4 C2f, C3k2, C2fCIB 3
S5 C2f, C3k2, C2fCIB, A2C2f 4
Operator (neck) N1–N4 C2f, C3k2, C2fCIB, A2C2f 4
@@
Each architecture configuration is converted into a 25-dimensional feature vector. Channel widths and block depths are represented as scalar values, while operator choices are one-hot encoded, allowing the tree-based predictor to split on each operator independently without imposing artificial ordering.
@@ 3.2 Self-Evolving Predictor
Evolutionary expansion (single round). Within each latency bucket, a target latency is sampled uniformly. An evolutionary algorithm (EA) [19] is then run with predicted mAP as the fitness function and real measured latency as the constraint. The EA runs for 100 generations with a population of 50, selecting the top 25% as parents, generating offspring via crossover (50%) or mutation (50%, rate 0.2), and carrying forward the top 10% as elite. The final top 5 architectures from each bucket are selected as candidates, yielding 50 new architectures per round.
Iterative refinement. Each batch of 50 new architectures is fully trained on COCO-mini under the same protocol, their configuration and ground-truth mAP are recorded, and the pairs are merged into the architecture pool. The predictor is then retrained on the expanded pool. This loop repeats for 10 rounds, growing the pool from 1,000 to 1,500 architectures. Crucially, each round’s newly added architectures are biased toward high-performance regions identified by the current predictor, progressively narrowing the distribution gap between predictor training and actual NAS search.
@@ 4.2 Main Results
Tab. 3 summarizes the surrogate predictor’s accuracy before and after Self-Evolving. The 10-model LightGBM ensemble trained on the initial 1,000 architectures achieves a validation sKT of 0.694. After 10 rounds of Self-Evolving, both $R^{2}$ and sKT improve: $R^{2}$ rises from 0.770 to 0.815 (+4.5%), and sKT from 0.694 to 0.752 (+5.8%). An sKT of 0.752 indicates strong ranking consistency, while the $R^{2}$ of 0.815 suggests that the predictor captures the majority of variance in architecture performance. These results show that our surrogate predictor is a high-fidelity proxy for the true performance landscape, and confirm that the Self-Evolving mechanism effectively sharpens the predictor’s ranking ability. Fig. 4 further visualizes the close agreement between predicted and actual mAP across all 1,500 architectures.
@@
Small (S) Medium (M) Large (L) Extra-Large (X)
@@
YOLO12 27.2 20.40 30.4 22.17 30.8 42.51 32.9 54.13
YOLO11 27.7 15.30 30.8 20.03 32.1 29.84 32.8 37.73
YOLOv10 25.8 15.89 29.7 20.01 30.1 25.83 30.5 36.84
YOLOv9 27.1 19.25 31.6 25.59 31.6 26.85 32.3 53.07
YOLOv8 26.2 10.95 29.4 17.21 30.8 23.30 32.9 33.71
Predictor-Guided Search Results.
@@
As shown in Tab. 4, the four predictor-discovered architectures Pareto-dominate all official YOLO baselines (v8–v12) across the full latency spectrum, achieving consistently higher mAP at equal or lower latency. The advantage is most pronounced in the small-model regime, where Ours-s surpasses YOLO11s by +4.2% mAP at comparable latency, and remains substantial at the large end, where Ours-x exceeds YOLO12x in mAP while being $1.5{\times}$ faster. These results confirm that the predictor not only achieves high ranking correlation (sKT) on held-out data, but also exhibits strong discriminative power in the top-performance regime—architectures it deems promising indeed yield superior performance after full retraining.
@@ 4.3 Ablation Studies
To disentangle whether the predictor improvement stems from the Self-Evolving mechanism or merely from increasing the training pool size, we conduct an ablation: we train 200 randomly sampled architectures on COCO-mini, add them to the initial 1,000 architectures, and retrain the ensemble predictor. For fair comparison, we also truncate Self-Evolving when the pool reaches 1,200 architectures. As shown in Tab. 6, random expansion yields $R^{2}=0.776$ and sKT=0.701=0.701, while Self-Evolving achieves $R^{2}=0.798$ and sKT=0.738=0.738. The clear gap (+0.022 $R^{2}$, +0.037 sKT) confirms that the gain is attributable to the targeted enrichment of high-performance architectures by the Self-Evolving loop, rather than to the increase in data size alone.
@@
4.4 Limitations and Future Work
The current benchmark is built on COCO-mini (10% of COCO) and measures latency on a single GPU (NVIDIA P40). To ensure the authenticity of latency measurements across different platforms, we do not build a surrogate latency predictor as we do for mAP; thus, when using this benchmark for NAS, one still needs to measure the latency of each architecture empirically. But this cost is acceptable compared to the error that a latency predictor would introduce on unseen hardware. Extending YOLO-NAS-Bench to the full COCO dataset, diverse hardware platforms (edge GPUs, mobile NPUs), and additional tasks such as instance segmentation and pose estimation are natural next steps.
@@ 5 Conclusion
We present YOLO-NAS-Bench, the first surrogate benchmark tailored to YOLO-style detectors. By designing a search space that spans channel width, block depth, and operator type across backbone and neck, and by training 1,000 architectures sampled via complementary strategies on COCO-mini, we establish an initial performance database. Building on this, our Self-Evolving Predictor iteratively enriches the database with high-value architectures, raising the ensemble LightGBM predictor’s sKT from 0.694 to 0.752. Architectures discovered by predictor-guided evolutionary search surpass all official YOLOv8–YOLO12 baselines at comparable latency on COCO-mini, confirming the predictor’s practical utility.
6 Acknowledgments
This work was supported by National Natural Science Foundation of China under Grant 62176007.